OpenStudy (anonymous):

OK! One more so I can go to bed already! The attendance at the Saturday night baseball game was three times the attendance at Sunday's game. In all, 56,000 fans attended the games. How many fans were at each game?

OpenStudy (anonymous):

So take the first sentence again. let A be the sAturday attendence and U be the sUnday attendence.

OpenStudy (anonymous):

What is the equation given by the first sentence.

OpenStudy (anonymous):

"Saturday's attendence was three times Sunday's attendance"

OpenStudy (anonymous):

Right. So A=Saturday and U=Sunday Equation: U*3=A

OpenStudy (anonymous):

Yes. Now if the total attendance for both nights was 56000. What is the equation that represents this fact?

OpenStudy (anonymous):

U+A=56000

OpenStudy (anonymous):

Yep. Now take what you know about A and substitute it out in that equation.

OpenStudy (anonymous):

Ok. This is where Im stuck! All I know is that A is 3 times U.

OpenStudy (anonymous):

Right. So if A = 3U then doesn't A + U = 3U + U ?

OpenStudy (anonymous):

O M G ! I think I got it! It just clicked! So my equation would be 3U+U=4U I then take 4U and divide it by the total 56000 I get 14,000. U=14,000 fans. RIGHT?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

i would then take the 14000 and subtract that by the total 56000 and get the answer for A. A=42000 Is this it?

OpenStudy (anonymous):

Yep, you got it.

OpenStudy (anonymous):

Alternatively you could have multiplied U by 3 to get A.

OpenStudy (anonymous):

OH WHAT JOY! PolPak, U f**kin rock! Thanks so much for being my personal tutor tonight! I really do appreciate it! I think I got it now!

OpenStudy (anonymous):

I'm really glad =)

OpenStudy (anonymous):

Have a great and relaxing night! Thanks again! ^_^

OpenStudy (anonymous):

You too.